mhchen:

There are 2 boxes. Box A has 1 black ball and 1 white ball Box B has 2 black balls and 1 white ball You select a box at random, then select a ball at random from that box. What is the probability that box A was selected given that you got a white marble?

mhchen:

P(A | white) = P(A and white) / P(white) = (1/2)(1/2) / (1/2)(1/2)+(1/2)(1/3)

mhchen:

= (1/4) / ((1/4)+(1/6)) = (1/4) / (5/12)

mhchen:

=12 / 20 = 3/5

mhchen:

right?

mhchen:

\[x^{2} = y\]